# LRQ-DiT: Log-Rotation Post-Training Quantization of Diffusion Transformers for Text-to-Image Generation

### v1

@@ 4.1 Experimental Setting
Models, Datasets and Metrics. Following prior studies (Zhao et al. 2024; Li et al. 2024b), we evaluated performance across different DiT models: PixArt-$\Sigma$ (steps=20, CFG=4.5), PixArt-$\alpha$ (steps=20, CFG=4.5), FLUX.1-schnell (steps=4, CFG=0.), and FLUX.1-dev (steps=50, CFG=3.5). We selected 1024 prompts from COCO (Chen et al. 2015), MJHQ-30K (Li et al. 2024a), and the summarized Densely Captioned Images (sDCI) (Urbanek et al. 2024) respectively to evaluate the generalization capability of our method LRQ-DiT. We evaluated both the quality and similarity of images generated by the quantized model. Specifically, we utilize the Fréchet Inception Distance (FID, the lower the better) (Heusel et al. 2017) to measure the distributional distance between images generated by the quantized and FP16 models. Then, we employ Image Reward (IR, the higher the better) (Xu et al. 2023) to estimate human preference ratings for the generated images. We adopt Peak Signal Noise Ratio (PSNR, the higher the better) to measure the numerical similarity of the generated images. Structural Similarity Index Measure (SSIM, the higher the better) is used to measure the luminance, contrast, and structure similarity of the images generated by the quantized and FP16 models.
@@
COCO dataset MJHQ dataset sDCI dataset COCO dataset MJHQ dataset sDCI dataset
@@
Table 3: Performance of LRQ-DiT (FLUX.1) text-to-image generation on different datasets.
@@
TLQ ARS FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$
$\times$ $\times$ 245.37 -2.04 151.59 -1.46 149.14 -1.43 232.04 -1.86 124.88 -0.87 121.65 -0.82
@@
$\checkmark$ $\times$  99.50 -0.29  65.61  0.00  64.87  0.01  76.02  0.38  55.14  0.61  54.85  0.62
@@
Table 4: Ablation studies of TLQ and ARS under different quantization settings and datasets on PixArt-$\alpha$.
Implementation Details and Baselines.
We have aligned the experimental settings of ViDiT-Q (Zhao et al. 2024), which include employing per-channel static quantization for weights and per-token dynamic quantization for activations, along with the migration strength of $0.5$ to accomplish difficulty transfer from activations to weights. Besides, we retain full precision for certain sensitive layers with a small number of parameters. Our PTQ method utilizes 4-10 prompts to obtain the smoothing factor, metric $\mathbf{J}$, rotation matrix $\mathbf{R}$, and permutation matrix $\mathbf{P}$ during the calibration phase, without the need for retraining the model or fine-tuning parameters. Therefore, it does not introduce additional overhead during the inference phase. Compared to the performance gain brought by our method, the additional computational overhead is acceptable. We compare LRQ-DiT with methods SmoothQuant (Xiao et al. 2023), QuaRot (Ashkboos et al. 2024), and the current state-of-the-art PTQ method ViDiT-Q (Zhao et al. 2024) to validate the effectiveness of our method. To ensure fairness in the comparison of experimental results, we rerun these methods on both PixArt and FLUX models using A6000 and A100 GPUs, respectively.
4.2 Main Results
We present the evaluation results of our method and comparative methods in Tables 1, 2, and 3, as detailed below:
PixArt. We have conducted validation for the text-to-image generation tasks on the COCO, MJHQ, and sDCI datasets using both PixArt-$\Sigma$ and PixArt-$\alpha$, and evaluated their performance under different settings: W3A4, W3A6, W3A8, W4A4, W4A6, and W4A8. The experimental results in Table 1 and 2 demonstrate that text-to-image generation cannot be achieved solely through a naive smoothing scheme. QuaRot mitigates the impact of outliers in the model, enabling the generation of moderately satisfactory images at high-bit width. However, once quantized to low-bit width, the model’s performance experiences a sharp decline. ViDiT-Q, as the current state-of-the-art PTQ method, performs well under 4-bit weight quantization, but the performance deteriorates once the weights are quantized to 3-bit. Our method achieves excellent performance levels under 3-bit weight quantization and further enhances the model’s performance under other quantization settings.
FLUX. Based on previous research findings (Li et al. 2024b) and our experimental observations, we have reached the following conclusion: due to FLUX inherent advantages, the models are capable of generating high-quality images even after low-bit quantization. Specifically, only using naive smoothing scheme to achieve 4-bit quantization, the model still exhibits great performance. We conducted 3-bit weight quantization experiments on the FLUX.1-schnell and FLUX.1-dev models and validated their performance on different datasets. The results in Table 3 demonstrate that our method enhances the model’s performance under the W3A4 quantization setting, with improvements observed in both the quality and similarity of the generated images.
Figure 3: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A6 Quantization.
@@ 4.3 Ablation Study
We conducted ablation experiments in Table 4 to validate the effectiveness of TLQ and ARS. Without TLQ and ARS, the model exhibits poor performance, and the generated images are also lacking in information. When ARS is applied independently to address the impact of mild and salient outliers in the model, the model’s performance is enhanced. When TLQ is applied independently to achieve more precise quantization of weights, the model’s performance also improves significantly. When TLQ and ARS are applied together, the quantized model achieves optimal performance in both image quality and similarity, providing strong evidence of the synergistic effect between TLQ and ARS.
@@ 4.4 Visualization Comparison
We generated images using QuaRot, ViDiT-Q and ours respectively on same prompts, with the specific results shown in Figure 3. LRQ-DiT demonstrates significant advantages in text-to-image generation. The foreground details in the generated images are presented with clarity and specificity, the depiction of motion scenes is precise and vivid, and at the same time, character recognition becomes much easier.
@@ 5 Conclusion
This paper introduces an efficient and accurate post-training quantization method LRQ-DiT for DiT models. LRQ-DiT identifies two key issues in model quantization and proposes corresponding solutions to address them. Specifically, TLQ utilizes twin-log quantization to allocate more quantization intervals to dense regions in weights, thereby reducing quantization errors and achieving precise quantization. Then, ARS identifies mild and salient outliers in activations through the metric and suppresses the influence of different types of outliers using an adaptive rotation scheme. We conduct quantization experiments with different bit settings on PixArt and FLUX models, and then validate the performance of text-to-image generation tasks on the COCO, MJHQ, and sDCI datasets. The experimental results demonstrate that LRQ-DiT not only enables the model to achieve satisfactory performance under 3-bit weight quantization but also enhances the model’s performance under other quantization settings, thereby exhibiting significant advantages.
Appendix A Visual Comparison
[…]
We provide more visual comparison results, as follows:
  • Figure 4 shows PixArt-$\Sigma$ under W3A4 quantization.
  • Figure 5 shows PixArt-$\Sigma$ under W3A6 Quantization.
  • Figure 6 shows PixArt-$\Sigma$ under W3A8 Quantization.
  • Figure 7 shows PixArt-$\alpha$ under W3A4 Quantization.
  • Figure 8 shows PixArt-$\alpha$ under W3A6 Quantization.
  • Figure 9 shows PixArt-$\alpha$ under W3A8 Quantization.
Figure 4: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A4 Quantization.
Figure 5: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A6 Quantization.
Figure 6: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A8 Quantization.
Figure 7: Visual Comparison of Generated Images from PixArt-$\alpha$ under W3A4 Quantization.
Figure 8: Visual Comparison of Generated Images from PixArt-$\alpha$ under W3A6 Quantization.
Figure 9: Visual Comparison of Generated Images from PixArt-$\alpha$ under W3A8 Quantization.

### v2

@@ 5.1 Experimental Setting
Models and Prompt sets
Following prior studies [62, 20], we evaluated performance across different DiT models: PixArt-$\Sigma$ (steps=20, CFG=4.5), PixArt-$\alpha$ (steps=20, CFG=4.5), FLUX.1-schnell (steps=4, CFG=0.), and FLUX.1-dev (steps=50, CFG=3.5). We selected 1024 prompts from COCO [7], MJHQ-30K [19], and the summarized Densely Captioned Images (sDCI) [40] respectively to evaluate the performance of LRQ-DiT in text-to-image generation tasks. In addition, we conducted performance validation on the text-to-video generation task on the OpenSORA v1.2 model, generating 51 frames and 240p videos. Specifically, we use the OpenSORA prompt set (including 10 prompts) and the Vbench prompt set (including 251 prompts) for verification.
Table 3: Performance of LRQ-DiT (FLUX.1) text-to-image generation on different prompt sets.
@@
COCO prompt set MJHQ prompt set sDCI prompt set COCO prompt set MJHQ prompt set sDCI prompt set
@@
Evaluation Metrics
We evaluated both the quality and similarity of images generated by the quantized model. Specifically, we utilize the Fréchet Inception Distance (FID, the lower the better) [13] to measure the distributional distance between images generated by the quantized and FP16 models. Then, we employ Image Reward (IR, the higher the better) [50] to estimate human preference ratings for the generated images. We adopt Peak Signal Noise Ratio (PSNR, the higher the better) to measure the numerical similarity of the generated images. Structural Similarity Index Measure (SSIM, the higher the better) is used to measure the luminance, contrast, and structure similarity of the images generated by the quantized and FP16 models. In addition, we use Imaging Quality and Aesthetic Quality to measure the quality of the video. Measure the consistency of generated videos using Background Consist, Subject Consist, Scene Consist, and Overall Consist. We also consider metrics such as PSNR, SSIM, VQA-Overall, Motion Smooth, Dynamic Degree, etc. for these generated videos.
Implementation Details and Baselines
We have aligned the experimental settings of ViDiT-Q [62], which include employing per-channel static quantization for weights and per-token dynamic quantization for activations, along with the migration strength of $0.5$ to accomplish difficulty transfer from activations to weights. Besides, we retain full precision for certain sensitive layers with a small number of parameters. Our PTQ method utilizes 4-10 prompts to obtain the smoothing factor, metric $\mathbf{J}$, rotation matrix $\mathbf{R}$, and permutation matrix $\mathbf{P}$ during the calibration phase, without the need for retraining the model or fine-tuning parameters. Therefore, it does not introduce additional overhead during the inference phase. Compared to the performance gain brought by our method, the additional computational overhead is acceptable. We compare LRQ-DiT with methods SmoothQuant [45], QuaRot [1], DuQuant [25] and the current state-of-the-art PTQ method ViDiT-Q [62] to validate the effectiveness of our method. To ensure fairness in the comparison of experimental results, we rerun these methods on both PixArt, FLUX and OpenSORA models using A6000 and A100 GPUs, respectively.
5.2 Main Results
PixArt
We have conducted validation for the text-to-image generation tasks on the COCO, MJHQ, and sDCI datasets using both PixArt-$\Sigma$ and PixArt-$\alpha$, and evaluated their performance under different settings: W3A4, W3A6, W3A8, W4A4, W4A6, and W4A8. The experimental results in Table 1 and 2 demonstrate that text-to-image generation cannot be achieved solely through a naive smoothing scheme. QuaRot mitigates the impact of outliers in the model, enabling the generation of moderately satisfactory images at high-bit width. However, once quantized to low-bit width, the model’s performance experiences a sharp decline. ViDiT-Q, as the current state-of-the-art PTQ method, performs well under 4-bit weight quantization, but the performance deteriorates once the weights are quantized to 3-bit. LRQ-DiT achieves excellent performance levels under 3-bit weight quantization and further enhances the model’s performance under other settings.
Table 4: Performance of the OpenSORA model in Text-to-Video Generation (51 frames, 240p) on the OpenSORA and VBench prompt sets.
Precison Method OpenSORA Set (10 prompts) Vbench Set (251 prompts)
PSNR $\uparrow$
SSIM $\uparrow$
VQA
Overall $\uparrow$
FVD
-FP16 $\downarrow$
Imaging
Quality $\uparrow$
Aesthetic
Quality $\uparrow$
Motion
Smooth $\uparrow$
Dynamic
Degree $\uparrow$
Background
Consist $\uparrow$
Subject
Consist $\uparrow$
Scene
Consist $\uparrow$
Overall
Consist $\uparrow$
W3A4 Smooth*  7.25 0.13 10.25 16.16 48.52 28.81 79.28 - 90.14 73.38 -  4.03
Quarot* 17.11 0.59 24.92  0.88 45.83 41.51 97.74 34.72 94.14 88.19 15.55 22.80
DuQuant* 17.27 0.59 21.09  1.38 43.31 40.51 98.32 29.16 94.63 88.06 20.05 21.03
ViDiT-Q* 17.52 0.58 21.07  1.42 44.58 42.36 98.08 33.33 93.87 89.08 20.56 22.79
Ours 17.87 0.63 30.29  1.45 49.43 47.77 98.56 40.27 95.13 93.00 30.08 25.52
W3A6 Smooth* 14.35 0.51 16.48 0.91 36.10 32.93 98.47 4.16 96.66 89.89 5.74 14.31
Quarot* 17.67 0.63 33.77 0.43 50.08 45.69 98.64 38.88 95.77 92.53 26.30 23.93
DuQuant* 17.00 0.58 23.74 0.80 50.71 45.45 98.71 43.05 95.92 92.24 32.55 23.75
ViDiT-Q* 17.52 0.61 31.02 0.79 49.70 46.89 98.53 43.05 95.43 92.10 25.79 24.17
Ours 18.40 0.67 41.48 0.32 52.61 51.28 98.93 52.78 96.37 95.31 32.99 25.85
W3A8 Smooth* 14.48 0.52 17.32 1.00 38.57 34.84 98.45 11.11 96.85 90.19 10.10 15.61
Quarot* 17.69 0.63 34.51 0.39 50.31 45.83 98.65 40.27 95.80 92.71 28.19 24.16
DuQuant* 17.45 0.62 35.87 1.11 49.30 46.82 98.73 37.50 95.82 92.92 32.04 24.27
ViDiT-Q* 17.49 0.61 33.28 1.07 49.75 47.22 98.57 45.83 95.42 92.53 29.21 24.20
Ours 18.63 0.67 44.17 0.30 52.69 52.00 98.92 48.62 96.40 95.51 33.22 25.75
W4A4 Smooth*  9.35 0.18  8.96 19.44 50.20 29.91 79.43 - 92.71 82.00 -  4.46
Quarot* 19.18 0.65 28.79  0.86 50.93 47.53 97.96 47.22 95.44 92.30 38.22 25.64
DuQuant* 20.29 0.68 28.90  0.94 53.08 47.90 98.30 51.38 95.38 92.55 34.01 25.69
ViDiT-Q* 20.19 0.68 28.50  0.92 51.60 47.30 98.14 54.16 95.27 92.01 35.61 25.84
Ours 20.33 0.69 28.83  0.71 51.79 47.95 98.34 55.56 95.56 92.87 37.80 25.87
W4A6 Smooth* 18.33 0.68 36.74 0.31 52.97 47.52 98.42 45.83 96.11 94.07 32.63 24.47
Quarot* 20.14 0.72 41.92 0.31 54.54 51.34 98.62 56.94 96.63 95.13 38.88 25.77
DuQuant* 20.47 0.72 40.69 0.39 55.45 50.71 98.66 50.00 96.46 94.86 36.33 26.14
ViDiT-Q* 21.61 0.75 38.90 0.40 55.57 50.89 98.66 52.77 96.52 94.83 36.77 25.98
Ours 22.78 0.79 42.04 0.44 56.57 52.25 98.85 48.62 96.90 95.28 41.28 26.68
FLUX
Based on previous research findings [20] and our experimental observations, we have reached the following conclusion: due to FLUX inherent advantages, the models are capable of generating high-quality images even after low-bit quantization. Specifically, only using naive smoothing scheme to achieve 4-bit quantization, the model still exhibits great performance. We conducted 3-bit weight quantization experiments on the FLUX.1-schnell and FLUX.1-dev models and validated their performance on different datasets. The results in Table 3 demonstrate that our method enhances the model’s performance under the W3A4 quantization setting, with improvements observed in both the quality and similarity of the generated images.
OpenSORA
We validate the performance of the OpenSORA v1.2 model in text-to-video generation tasks on 10 prompts in the OpenSORA prompt set and 251 prompts in the Vbench prompt set. We measure the quality, consistency, smoothness, dynamics, and other characteristics of the generated videos through various evaluation metrics. As shown in the Table 4, our method exhibits strong performance advantages compared to other comparative methods (SmoothQuant, QuaRot, DuQuant, and ViDiT-Q) in different bit settings and various prompt sets. Especially under low-bit quantization settings (W3A4, W3A6, and W3A8), LRQ-DiT exhibits stronger robustness and improves the performance of quantized models.
Table 5: Ablation studies of TLQ and ARS under different quantization settings and prompt sets on PixArt-$\alpha$.
@@ 5.2 Main Results
TLQ Hadamard DuQuant ARS FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$
$\times$ $\checkmark$ 245.37 -2.04 151.59 -1.46 149.14 -1.43 232.04 -1.86 124.88 -0.87 121.65 -0.82
$\times$ $\checkmark$ 269.46 -2.12 162.18 -1.49 166.63 -1.56 194.97 -1.60 165.78 -1.54 126.23 -1.01
@@
$\checkmark$ $\checkmark$  99.50 -0.30  65.61  0.01  64.87  0.01  76.02  0.38  55.14  0.61  54.85  0.62
$\checkmark$ $\checkmark$  99.47 -0.29  62.08  0.18  66.86 -0.19  76.10  0.15  60.39  0.47  57.53  0.66
@@
Figure 3: Visual Comparison of Images generated by PixArt-$\Sigma$ and Videos generated by OpenSORA using several different W3A6 quantization methods.
@@ 5.3 Ablation Study
We conducted ablation experiments in Table 5 to validate the effectiveness of TLQ and ARS. Without TLQ and ARS, the model exhibits poor performance, and images are also lacking in information. When ARS is applied independently to address the impact of mild and salient outliers in the model, the model’s performance is enhanced. When TLQ is applied independently to achieve more precise quantization of weights, the model’s performance also improves significantly. When TLQ and ARS are applied together, the quantized model achieves optimal performance in both image quality and similarity, providing strong evidence of the synergistic effect between TLQ and ARS. Compared to directly applying “Hadamard" or “DuQuant" rotation to all activations, ARS demonstrates significant advantages: it not only offers higher precision but also has quite limited computational overhead.
@@ 5.4 Visualization Comparison
We generated images using QuaRot, ViDiT-Q, and our method LRQ-DiT respectively on a set of prompts. The images and videos generated by different methods are shown in the Figure 3. After comparison, it can be seen that our method demonstrates significant advantages in generation tasks. The foreground details in the generated images by our method are presented with clarity and specificity, the depiction of motion scenes is precise and vivid, and at the same time, character recognition becomes much easier. Compared with other methods, the videos generated by our method have better quality, smoother flow, and stronger consistency.
Table 6: Performance Comparison between FP16 and INT8 GEMM (Triton Kernel).
Matrix Shape Block Shape Bit Time/ms TFLOPs SpeedUp
(2048,2048,2048) (64,64,64) FP16  0.290  59.283 $1.00\times$}
INT8  0.135 127.705 $\mathbf{2.15\times}$
(4096,4096,4096) (128,128,64) FP16  2.102  65.380 $1.00\times$}
INT8  0.858 160.164 $\mathbf{2.45\times}$
(8192,8192,8192) (64,256,32) FP16 16.810  65.409 $1.00\times$}
INT8  7.559 145.466 $\mathbf{2.22\times}$
5.5 SpeedUp and Latency
To validate the hardware efficiency of our method LRQ-DiT, we implemented a custom Triton-based kernel for INT8 matrix multiplication. In our setup, both weights and activations are quantized to INT8, and the matrix multiplication is performed entirely in INT8 before accumulation into higher-precision registers. All experiments were conducted on a single NVIDIA RTX 4090 GPU, where for each matrix shape, we automatically selected the optimal block configuration through heuristic search to maximize hardware utilization. Each kernel was executed for 10 iterations with warm-up runs to eliminate startup overhead, and the averaged runtime was reported. The results, summarized in Table LABEL:tab:lantency, demonstrate that INT8 execution achieves a consistent $2.15-2.45\times$ speedup over FP16 across different matrix sizes, with throughput improvements from approximately 65 TFLOPs to over 160 TFLOPs. These findings provide strong empirical evidence that the proposed quantization method not only preserves accuracy but also unlocks substantial acceleration on modern GPUs.
@@ 6 Conclusion
This paper introduces an efficient and accurate post-training quantization method LRQ-DiT for DiT models. LRQ-DiT identifies two key issues in model quantization and proposes corresponding solutions to address them. Specifically, Twin-Log Quantization (TLQ) utilizes log-based quantizer to allocate more quantization intervals to dense regions in weights, thereby reducing quantization errors and achieving precise quantization. Then, Adaptive Rotation Scheme (ARS) identifies mild and salient outliers in activations through the metric and suppresses the influence of different types of outliers using different rotation transformation. We conduct detailed experiments with various bit settings on PixArt and FLUX, and then validate the performance of text-to-image generation tasks on the COCO, MJHQ, and sDCI datasets. Besides, we also validate the performance of text-to-video generation on different prompt sets using the OpenSORA. The experimental results demonstrate that LRQ-DiT further improves the performance of the quantized models on different bit settings and prompt sets, especially in low-bit scenarios.